OpenStudy (anonymous):

find the derivative of y=lnx/x^2+1

OpenStudy (ash2326):

Is this your question? \[y=\frac{\ln x}{x^2+1}\]

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

use division rule here \[\large d(\frac uv)=\frac{vdu-udv}{v^2}\] \[\large\frac{dy}{dx}=\frac{(x^2+1)\times \frac{d}{dx}(\ln x)-\ln x\times \frac{d}{dx}(x^2+1)}{(x^2+1)^2}\]

OpenStudy (ash2326):

Can you do the rest?

OpenStudy (anonymous):

i get stuck :(